OpenStudy (ellaella02):

What is the x-intercept of the line with the equation 8x-1/3y=16 ? ( , 0)

OpenStudy (ellaella02):

@Directrix

OpenStudy (freckles):

set y to 0 and solve for x

OpenStudy (ellaella02):

um ok

OpenStudy (freckles):

\[8x-\frac{1}{3}y=16 \\ \\ 8x-\frac{1}{3}(0)=16 \text{ solve for } x \]

OpenStudy (anonymous):

8x-1/3y=16 8x-1/3(0)=16 8x-0=16 8x=16 8x/8=16/8 x=2 (2,0)

OpenStudy (freckles):

ok nevermind someone else can do all the work :p

OpenStudy (ellaella02):

Thank you so much! Do you think you could help me with three more? @Kobe95

OpenStudy (anonymous):

ill try my best what do you need help with?

OpenStudy (anonymous):

@EllaElla02

OpenStudy (ellaella02):

Which line is a graph of the equation: 2x + 5y = –10? Number graph ranging from negative five to five on the x axis and negative six to two on the y axis. Four lines are drawn in blue on the graph. Lines a and c have a positive slope and are parallel to each other. Lines b and d have a negative slope and are parallel to each other. Lines a and b intersect at (zero, two), and lines c and d intersect at (zero, negative two). A. line a B. line b C. line c D. line d

OpenStudy (ellaella02):

@Kobe95

OpenStudy (anonymous):

the line crossing through (0,-2) and (-5,0)

OpenStudy (anonymous):

those are your x and y intercepts

OpenStudy (anonymous):

in the picture i cant see the names of the lines

OpenStudy (ellaella02):

Okay one second

OpenStudy (ellaella02):

Number graph ranging from negative five to five on the x axis and negative six to two on the y axis. Four lines are drawn in blue on the graph. Lines a and c have a positive slope and are parallel to each other. Lines b and d have a negative slope and are parallel to each other. Lines a and b intersect at (zero, two), and lines c and d intersect at (zero, negative two).

OpenStudy (ellaella02):

I think it's line A.

OpenStudy (anonymous):

it is one that has a negative slope coming through the negative 5

OpenStudy (ellaella02):

When I zoomed in, Line A had the same points.

OpenStudy (anonymous):

its line D

OpenStudy (anonymous):

i zoomed in.

OpenStudy (ellaella02):

Oh um no I don't think so I really don't know because the more I look, he more I notice how they are all the same.

OpenStudy (ellaella02):

It is? Okay

OpenStudy (ellaella02):

Thank you (:

OpenStudy (anonymous):

was i right?

OpenStudy (ellaella02):

I have two more still.

OpenStudy (ellaella02):

I'll tell you when I finish the test (:

OpenStudy (anonymous):

ok

OpenStudy (ellaella02):

Can you help me with them?

OpenStudy (anonymous):

yes i will try

OpenStudy (ellaella02):

Which point is collinear with points B and C? A. (0, 0) B. (1, 1) C. (1, –5) D. (6, –8)

OpenStudy (anonymous):

my class will be over in like 6 minutes

OpenStudy (anonymous):

i believe c. (1,-5)

OpenStudy (ellaella02):

Okay, thank you (: I have one more

OpenStudy (anonymous):

ok

OpenStudy (ellaella02):

Complete the standard form of the equation –5y = 10 – 2x. 2 x + -5 y = 10

OpenStudy (ellaella02):

2x+-5y=10

OpenStudy (ellaella02):

I'm pretty sure that's the answer

OpenStudy (anonymous):

i believe so too.

OpenStudy (anonymous):

im not 100% but i do agree

OpenStudy (ellaella02):

YES I GOT AN 80% I PASSED YAAAAAAAAAASSSSSSSSSSSSSSSSSS

OpenStudy (ellaella02):

Number five ended up being A. 0,0

OpenStudy (ellaella02):

THANK YOU SO SO MUCH